# Multi-MeV electrons observed by CRRES during solar cycle 22

### v1

@@ Abstract
The Combined Release and Radiation Effects Satellite (CRRES) observed the response of the Van Allen radiation belts to peak solar activity within solar cycle 22. This study analyses the occurrence and loss timescales of relativistic electrons within the CRRES High Energy Electron Fluxometer (HEEF) dataset, including during several large geomagnetic storms that flooded the slot region with multi-MeV electrons and which allow the first definitive multi-MeV lifetimes to be calculated in this region. The HEEF loss timescales are otherwise broadly in agreement with those from later solar cycles but differences include longer-lasting sub-MeV electrons near the inner region of the outer belt and faster decaying multi-MeV electrons near geosynchronous orbit. These differences are associated with higher levels of geomagnetic activity, a phenomenon that enables the spread in the results to be parameterised accordingly. The timescales generally appear well-bounded by Kp-dependent theoretical predictions but the variability within the spread is however not always well-ordered by geomagnetic activity. This reveals the limits of pitch-angle diffusion in accounting for the decay of elevated electron fluxes following geomagnetic storms.
@@ 1 Introduction
The Van Allen electron radiation belts consist of two concentric tori encircling the Earth with energies extending up to several MeV [van Allen, 1959]. Satellites are designed to mitigate or avoid the impacts of these “killer” electrons which can lead to the triggering of phantom commands and component failure [e.g. Baker et al., 2013, Hands et al., 2018]. The precipitation of energetic particles also directly influences the ionisation and conductivities of the upper atmosphere, and its coupling back to the geospace environment [e.g. Roble and Rees, 1977, Orsolini et al., 2005]. The dynamics of trapped relativistic electron populations are thus a primary focus of space weather research.
@@
The electron radiation belts typically exhibit a two-belt structure, as commonly ordered by the Roederer [1970] L* parameter. A stable so-called inner belt peaks between L* $\approx$ 1–2 [Selesnick et al., 2007] and a highly variable outer belt extends between L* $\approx$ 3–8 [Glauert et al., 2018]. This canonical picture can however rapidly change during geomagnetic storms, with additional peaks in relativistic particles fluxes, i.e. new radiation belts, forming at various distances to the Earth, including inside the slot region. The first famous example occurred on 24 March 1991, where a large interplanetary shock near-instantaneously injected a new ultra-relativistic belt extending up to over 50 MeV inside the slot region [Blake et al., 1992]. A further significant event occurred during solar cycle 23 following the Hallowe’en storms of 2003 where the outer belt reformed inside the slot region and into the inner belt as a result of enhanced convection and erosion of the plasmasphere [Baker et al., 2004]. This again occurred in solar cycle 25, when the Mother’s/Gannon Day storm of May 2024 resulted in elevated fluxes forming inside the slot region [Pierrard et al., 2024]. The reordering of the outer belt into multiple distinct outer belts is also commonly occurring following geomagnetic storms [e.g. Pinto et al., 2018, Vampola, 1971, Baker et al., 2013, Chen et al., 2024].
@@
Elevated electron fluxes in the inner magnetosphere, following geomagnetic storms, are observed to decay exponentially across timescales of days-weeks, a phenomenon associated with pitch-angle scattering into the atmospheric loss cone [Meredith et al., 2006, 2009, Baker et al., 2004, Claudepierre et al., 2020]. This process provides a direct method to test diffusion theory [Lyons, 1974, Kennel et al., 1985, Glauert and Horne, 2005, Albert, 2005] and its application within 3-d radiation belt models [e.g. Fok et al., 2001, Shprits et al., 2009, Su et al., 2010, Reeves et al., 2012, Glauert et al., 2014]. As a general overview the following very low frequency (VLF) waves have found to be significant. Within the slot region, the dominant modes have been found to be plasmaspheric hiss [Lyons et al., 1972, Meredith et al., 2007], lightning-generated whistlers [e.g. Lauben et al., 2001] and magnetosonic waves [Ma et al., 2016, Wong et al., 2022]. Further out, hiss continues to play an important role for electron loss in the outer radiation belt both within the plasmasphere [Summers et al., 2004, Meredith et al., 2006] and in plasmaspheric plumes [Summers et al., 2008]. In addition, electromagnetic ion cyclotron (EMIC) [Miyoshi et al., 2008, Ross et al., 2021] and whistler-mode chorus waves [Shprits et al., 2007, Wang and Shprits, 2019, Wang et al., 2024] and outward radial diffusion to the magnetopause [Shprits et al., 2006, Mann et al., 2016] may also contribute to electron loss in the outer radiation belt. Anthopogenic VLF whistlers and Coulomb collisions have also been found to be effective in the inner belt and slot region [Ripoll et al., 2014, Cunningham et al., 2018, Ross et al., 2019b].
@@
Multi-MeV electron loss timescales have been empirically calculated within a few studies but the energy discrimination has often been limited to a single energy channel [e.g. West Jr. et al., 1981, Seki et al., 2005, Ripoll et al., 2015, Baker et al., 2007]. The Van Allen Probe electron lifetimes reported by Claudepierre et al. [2020] addressed this, determining decay rates within six differential energy channels extending from 1 – 4 MeV. This study addresses the analysis of multi-MeV electrons observed by CRRES during solar cycle 22, providing decay rates within six energy channels up to 4.55 MeV. Moreover, during the VAP mission multi-MeV electrons weren’t seen inside L*=2.8 [Baker et al., 2014] whereas in this study we find that CRRES regularly observed multi-MeV electrons inside the slot region. This enables the first calculation of well-resolved lifetimes in this region. To investigate the variability of the loss timescales we also parameterise the results by geomagnetic activity to allow an activity-dependent comparison of electron loss timescales with decay rates predicted by pitch angle diffusion theory.
@@
The paper is structured as follows. In Section 2 we outline the concept of pitch angle diffusion and the approach used to calculate the theoretical decay rates. In Section 3 we introduce the CRRES instrument and dataset and the automated algorithm used to derive the empirical loss timescales. In Section 4 we then present the lifetime analysis of multi-MeV electrons during the severe geomagnetic storms that flooded the slot region and the automated retrieval of lifetimes through the entire dataset. Section 5 then contrasts these with the theoretical solutions and previously reported loss timescales. Section 6 then summmarises the results of the study.
@@ 2 Pitch Angle Diffusion Theory
which can be solved as a boundary value problem [Albert, 1994, Meredith et al., 2006] for a given pitch angle distribution and as a function of chosen diffusion coefficients. The eigenvalues of the operator specify the time scales of the diffusion process [e.g. Albert and Shprits, 2009, Schulz and Lanzerotti, 1974]. The lowest order eigenvalues yield increasingly longer timescales and these appear dominant within observations reported so far at both equatorial and high latitudes [Meredith et al., 2006, Baker et al., 2007, Claudepierre et al., 2020]. However, this may not always be the case if the timescales of the diffusion process are longer than the timescales acting and the higher-order eigenvalues act across similar timescales to those examined [Albert, 1994, O’Brien et al., 2008]. The exponentially decaying solution in Equation 2, also allows empirical e-folding decays to be derived from observational data, as described in the next section, to serve as a direct comparison to these theoretical solutions.
@@ 3.1 Instrumentation
Table 1: Mid-point and ranges of the HEEF energy channels used in this study [Brautigm and Bell, 1995, McKellar, 1996].
@@
$<$0.80 $<$1.05 $<$1.70 $<$2.10 $<$2.50 $<$2.90 $<$3.30 $<$4.10 $<$4.95 $<$6.6 $<$8.55
In 2014, further empirical calibration factors were applied to the HEEF data to re-establish the utility of this dataset [Johnston et al., 2014]. Starting from a data set version with temperature and dead-time corrections applied, further corrections were implemented, including removal of data with incomplete pitch angle distributions and cross-calibration with the CRRES Medium Electron Analyser (MEA) [Vampola et al., 1992]. The resultant dataset consists of one-minute averages of differential electron fluxes separated into 5 pitch angle bins, with L* values derived from the Olson and Pfitzer [1977] magnetic field model. Table 1 outlines ten differential energy channels, eight of which are intrinsic to the instrument mode of operation and two additional differential channels (0.65 and 0.95 MeV) derived from differencing pairs of integral channels [Brautigm and Bell, 1995]. For further description of the cleaning of the data set and the contents, the reader is referred to Johnston et al. [2014].
@@ 3.2 Deriving Empirical Loss Timescales
Empirical electron lifetimes can be calculated using the same solution to Equation 1 through fitting the timescale, $\tau$, in Equation 2 directly to the HEEF data. We use a linear fit to the natural logarithm of the data to reduce interference from the largest fluxes. The derivation of electron lifetimes utilises daily averaged electron fluxes and their temporal evolution within individual energy bins over extended periods. The decays are fitted too using the exponential function over periods of over 25 data points ($\approx$ 5 days). The data is initially smoothed prior to fitting over a 11-point window ($\approx$1 days). To determine the quality of the fits we used the Pearson correlation coefficient, a value between -1 and 1, 0 being no correlation, -1 being a perfect negative correlation and 1 a perfect positive. Following periods of exponential decay, the flux levels exhibit significant variations likely representing a balance of production and loss and radial transport and the assumptions behind using Equations 2 are no longer valid.
This process is automated with predefined criteria specifying the determination of a fit to a decay timescale. This technique of identifying electron flux decays using an automated algorithm is analogous to that described in Meredith et al. [2009]. The data does not consist of daily averages, rather the average of each pass through 0.1 L shell ranges. The electron distribution functions, $A(\alpha_{0})$, are taken between local pitch angles 60-120 degrees and $\pm$15 degrees magnetic latitude. For automatic fitting, each flux data point is sequentially selected to be the starting point for a fit, with the initial end point determined by the given minimum fit range value. The resulting fit is used to find the Pearson coefficient and standard deviation. If a fit is found, i.e. the algorithm returns a negative Pearson coefficient with a value less than the threshold value of -0.95, the end point is incremented, with this process continuing until the Pearson value is no longer lower than the threshold. This incrementation of the fit length is continued within a given tolerance of Pearson values to avoid local minima. If the fit returns a Pearson with a value greater than the threshold, both the start and end points are incremented and the fitting process is repeated on the new data range until the Pearson value becomes less than the threshold, i.e. a new decay is found.
4 HEEF Dataset Analysis
@@
Figure 1 shows during the first half of the mission there were four strong storms and these are all associated with enhancements of 0.65 MeV electrons down to low values of L* with some events extending deep into the slot region and even the inner radiation belt. These storms tend also to be associated with flux enhancements at higher energies, up to 4.55 MeV. Interestingly, there is also a storm that doesn’t quite reach the threshold to be classified as a strong storm a few days before the second strong storm and this is also associated with enhancements of 0.65 MeV electrons to low L*. In sharp contrast there were 12 strong storms and two severe storms during the second half of the mission which result in a larger number of significant enhancements of 0.65 MeV electrons down to low L* and larger flux enhancements at higher energies.
@@
The Kp index notably shows different scalings to the Dst in that in the second half of the mission three storms reach Kp=9- (G4), an equivalence between them in contrast to the Dst index showing the single event on March 1991 reaching an intensity significantly greater than the others. The Kp index also reveals several peaks within each of the three G4 storms, for example in March 1991 showing two Kp=9- peaks corresponding to a first large compression reported early on 24 March [Blake et al., 1992] where the magnetopause was pushed inside geosychronous orbit for several hours [Elphic et al., 1991] and then a further large compression late on 24 March and 25 March, reported to drive large magnetopause oscillations near geosychronous orbit [Cahill and Winckler, 1992, Desai et al., 2021b]. The Kp index is a mid-latitude metric and this could be caused by motion of the auroral oval due to dynamic pressure enhancements associated with successive coronal mass ejections [Desai et al., 2020]. The Dst index is conversely an equatorial index, associated with southward directed interplanetary magnetic fields and build-up of the ring current, and this therefore captures the geomagnetic response across longer time-periods. Several studies have highlighted that radiation belt flux variations may be better correlated with the Kp index [Borovsky and Shprits, 2017, Wang et al., 2024], and the HEEF loss timescales and theoretical solutions in this study are parameterised as a function of the Kp index.
Figure 2 shows examples of the HEEF energy and pitch angle (PA) distributions corresponding to the data in Figure 1 during four periods of elevated fluxes and at four distinct radial distances, in the inner belt at L* $=1.5$, in the heart of the slot region at L* $=2.5$ and then in the outer belt at L* $=3.5$ and L* $=4.5$.
@@
Figure 4: Electron loss timescales in the slot region at L* = 2.5-2.6, as a function of energy for four time-periods during the CRRES mission noted in the legend.
@@
Figure 5: The 0.65 MeV electron energy number flux as a function of day number for the entire CRRES mission for (a-g) 1.6-1.7L*, 2.5-2.6L*, 3.0-3.1L*, 3.5-3.6L*, 4.0-4.1L*, 4.5-4.6L* and 5.0-5.1L*. The fitted exponential decays are overplotted in red
Figure 6: HEEF lifetimes of multi-MeV electrons presented for each energy bin as a function of L*. The raw data are coloured according to the average Kp index during the event, the black line shows the average fit with error bars showing the standard deviation of calculated lifetimes. Pitch angle diffusion theoretical predictions are shown at Kp = 0-1 and 4-5.
In Figures 1 and 2, an inner electron radiation belt is visible at L*$<$2 for the whole mission. In the lowest-energy channel 0, centred at 0.65 MeV, the dynamics of this inner belt correlate well with enhanced fluxes at higher L shells. The PA spectra at energy channel 0 (0.65 MeV) MeV at L* = 1.5 in Figure 2 also agrees well with expectations appearing symmetric about 90. The next energy channel 1 (0.95 MeV) is, however, devoid of fluxes in the inner belt. The higher energy channels do display counts here, but the non-uniform spectra are suggestive that this is proton contamination at multi-MeV energies in the inner zone. We are therefore unable to ascertain whether multi-MeV electrons appeared within the inner belt during the CRRES mission.
@@
The outer radiation belt appears highly variable throughout the CRRES mission driven by elevated geomagnetic activity. Following the geomagnetic storm of 24 March 1991, the multi-MeV electrons in the outer belt are separated into two further belts as examined by Kellerman et al. [2014] at lower energies. The corresponding energy and PA spectra are coherent up to 4.55 MeV and this multi-belt structure therefore also manifests at multi-MeV energies. The flux levels at around 10 counts appear highly variable, and we thus determine an approximate “noise floor" in the data below which we do not try to interpret signals as genuine. This effective noise floor, unfortunately, precludes the analysis in the study of the two highest energy channels 9 and 10, 5.75 and 7.5 MeV respectively, in this study.
@@
Within the slot region, see L*$=2.5$ panel in Figure 2 and Figures 1a-f, elevated flux levels start after 24 March 1991 (day 448) and extend right through to the end of the mission. The fluxes for the few orbits surrounding the injection of this new belt on day 448 [Blake et al., 1992] are not present in the dataset due to the cleaning procedures applied [Johnston et al., 2014] but the decay of this belt is visible following day 450. The corresponding energy and PA spectra in Figure 2j show coherent observations up to several MeV inside the slot region and throughout the outer belt. The further geomagnetic storms near days 525 and 560 in Figure 1 also produce elevated fluxes in the slot region of a similar or greater magnitude to those deriving from the 24 March 1991 event, with fluxes comparable to those in the outer belt. The Kp reaches similar magnitudes for these events, but the Dst index is significantly lower. These events are notable in that during solar cycle 24, multi-MeV electrons were unable to penetrate below L*=2.8 [Baker et al., 2014], with great solar activity in 2003 [Baker et al., 2004] and 2024 [Pierrard et al., 2024] required to overcome this. These also enable us to calculate slot-region loss timescales at multi-MeV energies, something that was not possible during the Van Allen probe era.
@@
To further examine the slot-filling events, Figure 3 shows the data within an L*=0.1 bin, centred at L* = 2.5, in each energy channel up to 4.55 MeV. Most of the energy channels exhibit little data in the slot region, but three decay periods following storms in the latter part of the mission are fitted to, overlaid as red lines, with Pearson correlation coefficients all lower than –0.95, together with a decay near day 240. Figure 4 shows the resultant lifetimes as a function of energy from these periods. In the first event we only observe an electron response at 0.65 MeV, whereas in the next three events all energies are affected. The trends show increased lifetimes at sub-MeV energies which decreases to a minimum near 1 MeV associated with increased pitch angle diffusion into the loss cone [Glauert et al., 2024, see Fig 3 therein]. The lifetimes then increase again and level out at multi-MeV energies.

### v2

# Multi-MeV electron occurrence and lifetimes in the outer radiation belt and slot region during the maximum of solar cycle 22

@@ Abstract
The Combined Release and Radiation Effects Satellite (CRRES) observed the response of the Van Allen radiation belts to peak solar activity within solar cycle 22. This study analyses relativistic and ultra-relativistic electron occurrence and loss timescales within the CRRES High Energy Electron Fluxometer (HEEF) dataset, including during several strong and severe geomagnetic storms that all, remarkably, flooded the slot region with multi-MeV electrons. These allow the first definitive multi-MeV electron lifetimes to be calculated in this region and indicate an elevated risk to satellites in slot region orbits during periods of heightened solar activity. The HEEF outer belt loss timescales are broadly in agreement with those from later solar cycles, but differences include longer-lasting sub-MeV electrons near the inner region of the outer belt and faster-decaying multi-MeV electrons near geosynchronous orbit. These differences are associated with higher levels of geomagnetic activity, a phenomenon that enables the spread in the results to be parameterised accordingly. The timescales generally appear well-bounded by Kp-dependent theoretical predictions, but the variability within the spread is not always well-ordered by geomagnetic activity. This suggests the limitations of using pitch-angle diffusion to account for the decay of elevated electrons following geomagnetic storms, and the need for more sophisticated space weather indices for radiation belt forecasting.
@@ 1 Introduction
[1] The Van Allen electron radiation belts consist of two concentric tori encircling the Earth with energies extending up to several MeV [van Allen, 1959]. Satellites are designed to mitigate or avoid the impacts of these “killer” electrons which can lead to the triggering of phantom commands and component failure [e.g. Baker et al., 2013, Hands et al., 2018]. The precipitation of energetic particles also directly influences the ionisation and conductivities of the upper atmosphere, and its coupling back to the geospace environment [e.g. Roble and Rees, 1977, Orsolini et al., 2005]. The dynamics of trapped relativistic electron populations are thus a primary focus of space weather research and forecasting efforts.
@@
[2] The electron radiation belts typically exhibit a two-belt structure, as commonly ordered by the Roederer [1970] L* parameter. A stable so-called inner belt peaks between L* $\approx$ 1–2 [Selesnick et al., 2007] and a highly variable outer belt extends between L* $\approx$ 3–8 [Glauert et al., 2018]. This canonical picture can however rapidly change during geomagnetic storms, with additional peaks in the relativistic particles flux, i.e. new radiation belts, forming at various distances to the Earth, including inside the slot region. The first well-known example occurred on 24 March 1991, where a large interplanetary shock near-instantaneously injected a new ultra-relativistic belt extending up to over 50 MeV inside the slot region [Blake et al., 1992]. Elevated energetic particles fluxes during this March 1991 event resulted in two satellite outages, and single event upsets (SEUs) and solar cell degradation on orbiting satellites [Mullen et al., 1991, Ishikawa et al., 2013, Shea and Smart, 1993]. A further significant event occurred during solar cycle 23 following the Hallowe’en storms of 2003 where the outer belt reformed inside the slot region and into the inner belt as a result of enhanced convection and erosion of the plasmasphere [Baker et al., 2004]. This similarly caused a satellite outage and extensive anomalies [Cho, 2005, NOAA Space Environment Center, 2004]. This again occurred in solar cycle 25, when the Mother’s/Gannon Day storm of May 2024 resulted in elevated fluxes forming inside the slot region [Pierrard et al., 2024, Li et al., 2025]. The reordering of the outer belt into multiple distinct outer belts is also commonly occurring following geomagnetic storms [e.g. Pinto et al., 2018, Vampola, 1971, Baker et al., 2013, Chen et al., 2024].
@@
[3] Elevated electron fluxes in the inner magnetosphere, following geomagnetic storms, are observed to decay exponentially across timescales of days-weeks, a phenomenon associated with pitch-angle scattering into the atmospheric loss cone [Meredith et al., 2006, 2009, Baker et al., 2004, Claudepierre et al., 2020]. This process provides a direct method to test diffusion theory [Lyons, 1974, Kennel et al., 1985, Glauert and Horne, 2005, Albert, 2005] and its application within 3-d radiation belt models [e.g. Fok et al., 2001, Shprits et al., 2009, Su et al., 2010, Reeves et al., 2012, Glauert et al., 2014]. As a general overview the following plasma waves have found to be significant. Within the slot region, the dominant modes have been found to be plasmaspheric hiss [Lyons et al., 1972, Meredith et al., 2007], lightning-generated whistlers [e.g. Lauben et al., 2001] and magnetosonic waves [Ma et al., 2016, Wong et al., 2022]. Further out, hiss continues to play an important role for electron loss in the outer radiation belt both within the plasmasphere [Summers et al., 2004, Meredith et al., 2006] and in plasmaspheric plumes [Summers et al., 2008]. In addition, electromagnetic ion cyclotron (EMIC) [Miyoshi et al., 2008, Ross et al., 2021] and whistler-mode chorus waves [Shprits et al., 2007, Wang and Shprits, 2019, Wang et al., 2024] and outward radial diffusion to the magnetopause [Shprits et al., 2006, Mann et al., 2016] may also contribute to electron loss in the outer radiation belt. Anthopogenic VLF whistlers and Coulomb collisions have also been found to be effective in the inner belt and slot region [Ripoll et al., 2014, Cunningham et al., 2018, Ross et al., 2019b].
@@
[4] Multi-MeV electron loss timescales have been empirically calculated within a few studies but the energy discrimination has often been limited to a single energy channel [e.g. West Jr. et al., 1981, Seki et al., 2005, Ripoll et al., 2015, Baker et al., 2007]. The Van Allen Probe electron lifetimes reported by Claudepierre et al. [2020] addressed this, determining decay rates within six differential energy channels extending from 1 – 4 MeV. This study provides analysis of multi-MeV electrons observed by CRRES during the maximum of solar cycle 22, providing decay rates within six energy channels up to 4.55 MeV. Moreover, during the VAP mission multi-MeV electrons weren’t seen inside L*=2.8 [Baker et al., 2014] whereas in this study we find that CRRES regularly observed multi-MeV electrons inside the slot region. This enables the first calculation of well-resolved lifetimes in the slot region. To investigate the variability of the loss timescales we also parameterise the results by geomagnetic activity to perform an activity-dependent comparison of empirical electron loss timescales with predictions by pitch angle diffusion theory.
@@
[5] The paper is structured as follows. In Section 2 we outline the concept of pitch angle diffusion and the approach used to calculate the theoretical decay rates. In Section 3 we introduce the CRRES instrument and dataset and the automated algorithm used to derive the empirical loss timescales. In Section 4 we then present the geomagnetic activity-dependent lifetime analysis of multi-MeV electrons during the severe geomagnetic storms that flooded the slot region and the automated retrieval of lifetimes through the entire dataset. Section 6 then contrasts these with the theoretical solutions and previously reported loss timescales. Section 7 then summmarises the results of the study.
@@ 2 Pitch Angle Diffusion Theory
a second order difference equation that can be solved as a boundary value problem [Albert, 1994, Meredith et al., 2006] for a given pitch angle distribution and as a function of chosen diffusion coefficients. The eigenvalues of the operator specify the time scales of the diffusion process [e.g. Albert and Shprits, 2009, Schulz and Lanzerotti, 1974] which Sturm-Liouville theory assumptions dictate are real and ordered [O’Brien et al., 2008, Claudepierre et al., 2022]. The lowest order eigenvalues therefore yield increasingly longer timescales and these have appeared dominant within observations reported so far at both equatorial and high latitudes in the outer radiation belt [Meredith et al., 2006, Baker et al., 2007, Claudepierre et al., 2020]. However, this may not always be the case if the timescales of the diffusion process are longer than the timescales acting and the higher-order eigenvalues act across similar timescales to those examined as in the inner slot region and inner radiation belt [Meredith et al., 2009, Broll et al., 2023]. The exponentially decaying solution in Equation 2, allows empirical e-folding decays to be derived from observational data, as described in the next section, to serve as a direct comparison to these theoretical solutions.
@@ 3.1 Instrumentation
Table 1: Mid-point and ranges of the HEEF energy channels used in this study [Brautigam and Bell, 1995, McKellar, 1996].
@@
<<0.80 <<1.05 <<1.70 <<2.10 <<2.50 <<2.90 <<3.30 <<4.10 <<4.95 <<6.6 <<8.55
In 2014, further empirical calibration factors were applied to the HEEF data to re-establish the utility of this dataset [Johnston et al., 2014b]. Starting from a data set version with temperature and dead-time corrections applied, further corrections were implemented, including removal of data with incomplete pitch angle distributions and cross-calibration with the CRRES Medium Electron Analyser (MEA) [Vampola et al., 1992]. The resultant dataset consists of one-minute averages of differential electron fluxes separated into 5 pitch angle bins, with L* values derived from the Olson and Pfitzer [1977] magnetic field model. Table 1 outlines ten differential energy channels, eight of which are intrinsic to the instrument mode of operation and two additional differential channels (0.65 and 0.95 MeV) derived from differencing pairs of integral channels [Brautigam and Bell, 1995]. For further description of the cleaning of the data set and the contents, the reader is referred to Johnston et al. [2014b].
@@ 3.2 Deriving Empirical Loss Timescales
Empirical electron lifetimes can be calculated using the same solution to Equation 1 through fitting the timescale, $\tau$, in Equation 2 directly to the HEEF data. We use a linear fit to the natural logarithm of the data to reduce interference from the largest fluxes. The derivation of electron lifetimes utilises daily averaged electron fluxes and their temporal evolution within individual energy bins over extended periods. The data does not consist of daily averages, rather the average of each pass through 0.1 L shell ranges. In the outer belt and slot region the electron distribution functions, $A(\alpha_{0})$, are summed between local pitch angles 60-120 degrees to maximise counts and taken for observations between $\pm$15 degrees magnetic latitude. In the inner belt, 90 degrees pitch angle fluxes are used only, as the lifetimes here are highly pitch angles dependent. The data is smoothed over a period of 1-2 days and decays fitted using the exponential function over periods not less than 5 days, although the results are relatively insensitive to these specific parameters. To determine the quality of the fits we used the Pearson correlation coefficient, a value between -1 and 1, 0 being no correlation, -1 being a perfect negative correlation and 1 a perfect positive, with a threshold of –0.95 for determining an accurate timescale. Following periods of exponential decay, the flux levels exhibit significant variations likely representing a balance of production and loss and radial transport and the assumptions behind using Equations 2 are no longer valid.
This process is automated with predefined criteria specifying the determination of a fit to a decay timescale. For automatic fitting, each flux data point is sequentially selected to be the starting point for a fit, with the initial end point determined by the given minimum fit range value. The resulting fit is used to find the Pearson coefficient and standard deviation. If a fit is found, i.e. the algorithm returns a negative Pearson coefficient with a value less than the threshold, the end point is incremented, with this process continuing until the Pearson value is no longer lower than the threshold. This incrementation of the fit length is continued within a given tolerance of Pearson values to avoid local minima. If the fit returns a Pearson with a value greater than the threshold, both the start and end points are incremented and the fitting process is repeated on the new data range until the Pearson value becomes less than the threshold, i.e. a new decay is found.
4 Geomagnetic Storms in the CRRES Era
@@
Figure 1 shows during the first half of the mission there were four strong storms and these are all associated with enhancements of 0.65 MeV electrons down to low values of L* with some events extending deep into the slot region and even the inner radiation belt. These storms tend also to be associated with flux enhancements at higher energies, up to 4.55 MeV. Interestingly, there is also a moderate storm that doesn’t quite reach the threshold to be classified as strong a few days before the second moderate storm and this is also associated with enhancements of 0.65 MeV electrons to low L*. In sharp contrast there were twelve strong storms and two severe storms during the second half of the mission which result in a larger number of significant enhancements of 0.65 MeV electrons down to low L* and larger flux enhancements at higher energies.
@@
The Kp index notably shows different scalings to the Dst in that in the second half of the mission three storms reach Kp=9- (G4), an equivalence between them in contrast to the Dst index showing the single event on March 1991 reaching an intensity significantly greater than the others. The Kp index also reveals several peaks within each of the three G4 storms, for example in March 1991 showing two Kp=9- peaks corresponding to a first large compression reported early on 24 March [Blake et al., 1992] where the magnetopause was pushed inside geosychronous orbit for over five hours [Elphic et al., 1991] and then a further large compression late on 24 March and 25 March, reported to drive large magnetopause oscillations near geosychronous orbit [Cahill and Winckler, 1992, Desai et al., 2021b]. The Kp index is a mid-latitude metric and this could be caused by motion of the auroral oval due to dynamic pressure enhancements associated with successive coronal mass ejections [Koehn et al., 2022]. The Dst index is conversely an equatorial index, associated with southward directed interplanetary magnetic fields and build-up of the ring current, and this therefore captures the geomagnetic response across longer time-periods. Several studies have highlighted that radiation belt flux variations may be better correlated with the Kp index [Borovsky and Shprits, 2017, Wang et al., 2024].
Significant attention has been devoted to deriving the properties of the interplanetary shock that struck on 24 March 1991 (day 447) due to the rapid nature that the slot region was filled with indications that it reached 1500 km/s at the Earth [Blake et al., 1992, Li et al., 1993, Hudson et al., 1997, Elkington et al., 2004]. The June storm (day 524) and subsequence disturbances persisted for nearly 20 days, and featured four geomagnetic storms in total of varying intensity, the magnetospheric responses identified as likely reflecting successive high-speed solar wind structures and periods of strong southward solar wind magnetic field [Gao et al., 1997]. Although the solar wind conditions have not been analysed, the July 1991 (day 554) storm was similarly inferred to derive from a large enhancement in dynamic pressure which pushed the magnetopause inside geosynchonous orbit for 3 hours, and large sudden storm commencement electric field that induced notable effects in the ionosphere [Burke et al., 2000, Wilson et al., 2001]. The storm of 04 June 1991 (day 520) received less attention, but is listed 57 in the most extreme storms between 1868 and 2010, as ordered by the AA index [Vennerstrom et al., 2016]. This is two places above the 24 March 1991 event which sits at 59 and the July event is not listed. This AA-ordering further highlights inconsistencies between different geomagnetic indices. The lack of continuous upstream solar wind observation in 1991, with only intermittent observations from IMP8 when orbiting in the solar wind, unfortunately precludes further analysis of the upstream drivers of these storms observed by CRRES.
@@
Figure 4: Electron loss timescales in the slot region at L* = 2.5-2.6, as a function of energy for four doy time-periods during the CRRES mission noted in the legend.
@@
Figure 5: The 0.65 MeV electron energy number flux as a function of day number for the entire CRRES mission for (a-g) 1.5-1.6L*, 2.5-2.6L*, 3.0-3.1L*, 3.5-3.6L*, 4.0-4.1L*, 4.5-4.6L* and 5.0-5.1L*. The fitted exponential decays are overplotted in red
@@
The outer radiation belt appears highly variable throughout the CRRES mission driven by elevated geomagnetic activity. Following the geomagnetic storm of 24 March 1991, the multi-MeV electrons in the outer belt are separated into two further belts as examined by Kellerman et al. [2014] at lower energies. The corresponding energy and PA spectra are coherent up to 4.55 MeV and this multi-belt structure therefore also manifests at multi-MeV energies. The flux levels at low counts appear highly variable, and we thus determine an approximate “noise floor" in the data below which we do not try to interpret signals as genuine. This high noise floor, unfortunately, precludes the analysis in the study of the two highest energy channels 9 and 10, 5.75 and 7.5 MeV respectively, in this study.
Figure 2 shows HEEF energy and pitch angle (PA) distributions corresponding to the data in Figure 1 during three large geomagnetic storms and a lesser fourth storm on 26 August 1991, day 238. For each of these, the spectra are displayed at four distinct radial distances, in the inner belt at L* =1.5=1.5, in the heart of the slot region at L* =2.5=2.5 and then in the outer belt at L* =3.5=3.5 and L* =4.5=4.5. In Figures 1 and 2, an inner electron radiation belt is visible at L*<<2 for the whole mission. In the lowest-energy channel 0, centred at 0.65 MeV, the dynamics of this inner belt correlate well with enhanced fluxes at higher L shells. The PA spectra at energy channel 0 (0.65 MeV) MeV at L* = 1.5 in Figure 2 also agrees well with expectations appearing symmetric about 90. The next energy channel 1 (0.95 MeV) is, however, devoid of fluxes in the inner belt. The higher energy channels do display counts here, but the non-uniform spectra are suggestive that this is proton contamination at multi-MeV energies in the inner zone. We are therefore unable to ascertain whether MeV and multi-MeV electrons appeared within the inner belt during the CRRES mission.
@@
Within the slot region, see L*=2.5=2.5 panel in Figure 2 and Figures 1a-f, elevated flux levels start after 24 March 1991 (day 448) and extend right through to the end of the mission. The fluxes for the few orbits surrounding the injection of this new belt on day 448 [Blake et al., 1992] are not present in the dataset due to the cleaning procedures applied [Johnston et al., 2014b] but the decay of this belt is visible following day 450. The corresponding energy and PA spectra in Figure 2j show coherent observations up to several MeV inside the slot region and throughout the outer belt. The further geomagnetic storms after days 524 and 554 in Figure 1 also produce elevated fluxes in the slot region of a similar or greater magnitude to those deriving from the 24 March 1991 event, with fluxes comparable to those in the outer belt. The extended June 1991 disturbances [Gao et al., 1997] correlates with an increase in the slot region electron flux over an extended time period, in contrast to the March 1991 more sudden increase. The July increase simillarly occurs over an extended period.
The Kp reaches similar magnitudes for these events, but the Dst index is significantly lower. These events are notable in that during solar cycle 24, multi-MeV electrons were unable to penetrate below L*=2.8 [Baker et al., 2014], with great solar activity in 2003 [Baker et al., 2004] and 2024 [Pierrard et al., 2024, Li et al., 2025] required to overcome this. These events enable us to calculate slot-region loss timescales at multi-MeV energies, something that was not possible during the Van Allen probe era.
@@
To examine the persistence of the multi-MeV electrons within these slot-filling events, Figure 3 shows the data within an L*=0.1 bin, centred at L* = 2.5, in each energy channel up to 4.55 MeV.The first storm only displays data at the lowest energy channels but three decay periods following storms in the latter part of the mission are fitted to at all energy channels, overlaid as red lines, with Pearson correlation coefficients all lower than –0.95. Figure 4 shows the resultant lifetimes as a function of energy from these periods. The trends show increased lifetimes at sub-MeV energies which decreases to a minimum near 1 MeV associated with increased pitch angle diffusion into the loss cone [Glauert et al., 2024, see Fig 3 therein]. The lifetimes then increase again and level out at multi-MeV energies.
Figure 6: HEEF outer belt and slot region lifetimes of multi-MeV electrons presented for each energy bin as a function of L*. The raw data are coloured according to the average Kp index during the decay period, the black line shows the average fit with error bars showing the standard deviation of calculated lifetimes. Pitch angle diffusion theoretical predictions are shown at Kp = 0-1 and 4-5.
5 HEEF Electron Lifetimes